DIFFERENTIAL GEOMETRY/PDE SEMINAR

Monday, May 3 Thomson 325 4:00 pm

Deforming real hypersurfaces by the trace of the Levi form

Gerhard Huisken (Princeton University)

The lecture describes joint work with W.Klingenberg concerning the deformation of real hypersurfaces of a Kähler manifold in direction of their real unit normal with speed determined by the trace of the Levi form. This evolution equation is a subparabolic system, for which short time existence and a number of qualitative properties will be demonstrated.